# Density Descent for Diversity Optimization

### v1

@@ 6.3 Analysis
Figure 3: Coverage and cross-entropy (CE) after 5,000 iterations of each algorithm in all domains. We report the mean over 10 trials, with error bars showing the standard error of the mean. Higher coverage and lower cross-entropy are better.
@@
We attribute the high coverage of DDS-KDE and DDS-CNF on these domains to the continuity of their density estimate, which prevents DDS-KDE from converging prematurely. For example, on LP, our algorithms discover more solutions near the edges of the feature space than CMA-MAE (Appendix G). The passive archive maintained by CMA-MAE converges as all the solutions fall in previously discovered cells (Fontaine et al.,, 2020). In contrast, the continuity of our density estimate and the online buffer updates facilitate DDS algorithms to achieve slow but continual progress in exploring the feature space (Fig. 3, LP and Arm Repertoire). This is because the shape of the continuous density estimate always changes slightly after each iteration of the algorithm.
For multi-feature LP, we observe that algorithms leveraging continuous representations of the feature space, i.e., DDS and NS, explore the feature space much faster than other algorithms driven by discrete feature space representations, e.g., CMA-MAE (Fig. 3). This is because CMA-MAE is optimizing on a centroidal Voronoi tessellation with 10,000 cells, where each cell has significantly more volume compared to that of a $100\times 100$ grid due to the increased dimensionality of the feature space. Hence, more solutions map to the same cells, making it more difficult to find solutions outside of previously explored cells. These results show that DO algorithms based on continuous density estimates scale better with the dimensionality of the feature space.
@@
We attribute the difference between the performance of DDS-KDE and DDS-CNF to the bandwidth parameter in KDE, which allows us to adjust the smoothness of the KDE (1(a)). On the other hand, CNF lacks explicit control over its smoothness. Thus, while we adjust the smoothness of the KDE to improve performance for DDS-KDE (Sec. 5), we can not apply the same techniques to boost the performance of DDS-CNF.
7 Limitations and Future Work
@@
We introduce two new diversity optimization algorithms, DDS-KDE and DDS-CNF. We provide theoretical insight into the connection between DDS-KDE and novelty search and show that both DDS algorithms excel at discovering diverse solutions.
@@
We envision several extensions to DDS. While we proposed two DDS variants, DDS is a general method that can integrate with a wide variety of distribution fitting techniques. DDS can be combined with parametric estimators like mixture models (McLachlan et al.,, 2019) and moment matching (Han et al.,, 2018), as well as with non-parametric estimators like vector quantization (Kohonen,, 1990) and generative models (Goodfellow et al.,, 2020). In general, we believe DDS will interface with any method that learns a continuous density representation of the feature space.
@@
For DDS-KDE, while we assumed a constant bandwidth, prior work has adapted the bandwidth online (Kristan et al.,, 2011) and varied the bandwidth based on query location (Terrell and Scott,, 1992). These methods could be applied to secure tighter uniform stability bounds for the KDE, making feature space exploration more efficient.
@@
A key limitation of our algorithms is their significant computational cost. Querying the KDE incurs a runtime of $O(|{\mathcal{B}}|)$, where $|{\mathcal{B}}|$ is the buffer size, and fine-tuning the CNF is computationally expensive due to backpropagation, even when accelerated on a GPU. These issues could be mitigated by accelerating KDE computation (O’Brien et al.,, 2016; Yang et al.,, 2003) and CNF training efficiency (Onken et al.,, 2021; Huang and Yeh,, 2021).
@@
Finally, we are excited to see how the underlying insight of DDS — leveraging continuous density representations of the feature space — can improve the exploration power of QD algorithms, especially in high-dimensional feature spaces. Applying DDS to QD will improve its performance on applications such as scenario generation for complex systems (Fontaine et al.,, 2021), generative design (Gaier et al.,, 2018), damage recovery in robotics (Cully et al.,, 2015), reinforcement learning (Batra et al.,, 2023), and procedural content generation (Gravina et al.,, 2019).
@@ Appendix A Hyperparameters
Following prior work (Fontaine et al.,, 2020), all algorithms run with 15 emitters — emitters (Tjanaka et al.,, 2023) are objects that generate solutions for QD algorithms. For MAP-Elites (line), the emitter applies the Iso+LineDD operator to random solutions from the archive (Vassiliades and Mouret,, 2018). For other algorithms, the emitters are individual instances of CMA-ES. In each experiment, we run the algorithm for 5000 iterations, where each emitter generates $\lambda=36$ solutions per iteration. Following prior work (Fontaine and Nikolaidis,, 2023), CMA-ES is configured with the $\mu$ selection rule, basic restart rule, and an initial solution ${\bm{x}}_{0}={\bm{0}}$.
For LP, Arm Repertoire, and Deceptive Maze, MAP-Elites-based algorithms (CMA-MAE, CMA-ME, and MAP-Elites (line)) run on a grid archive of $100\times 100$ following (Fontaine and Nikolaidis,, 2023). For Multi-feature LP, since maintaining a grid tessellation would require an exponential number of cells in the number of dimensions, we use a CVT archive with 10,000 cells. Due to randomness in the generation of the CVT cells, we pre-generate the centroids and use the same centroids for all algorithms.
Our DDS algorithms (DDS-KDE and DDS-CNF) maintain a buffer of capacity 10,000. DDS-KDE uses a Gaussian kernel with a bandwidth parameter of $h=25.6,10,0.01,2.56$ for Linear Projection, Arm Repertoire, Deceptive Maze, and Multi-feature LP, respectively. The bandwidth is found through empirical fine-tuning (Appendix B).
Table 1: Detailed hyperparameter specifications for each algorithm and domain.
DDS-KDE Initial step size $\sigma_{0}$ 1.5 0.5 1.5 1.5
Bandwidth $h$ 25.6 10 0.01 2.56
DDS-CNF Initial step size $\sigma_{0}$ 0.5 1.5 1.5 0.5
NS Initial step size $\sigma_{0}$ 0.5 0.5 0.5 0.5
Nearest neighbors $k$ 100 100 100 100
CMA-MAE Initial step size $\sigma_{0}$ 0.5 0.2 0.5 0.5
Threshold min 0 0 0 0
Learning rate $\alpha$ 0.01 0.01 0.01 0.01
CMA-ME Initial step size $\sigma_{0}$ 0.5 0.2 0.5 0.5
Line variance $\sigma_{\mathrm{line}}$ 0.2 0.2 0.2 0.2
@@ Appendix B Bandwidth Selection for DDS-KDE
The bandwidth parameter $h$ in KDE controls the smoothness of the estimation (1(a)). While prior work (Silverman,, 1986; Rudemo,, 1982; Bowman,, 1984) has developed bandwidth selection methods for accurate density estimations, the optimal bandwidth for DDS-KDE may require an over-smoothing of the KDE (Sec. 5).
To examine the effect of the bandwidth parameter on the performance of DDS-KDE, we perform ablation experiments on LP, Arm Repertoire, and Deceptive Maze. While applying the same hyperparameters as described in Appendix A, we vary the bandwidth $h\in(0,W]$, where $W$ is the bound of the feature space defined for each domain (Sec. 6.2). For instance, in the LP domain, $W_{\mathrm{LP}}=5.12\cdot\frac{n}{2}$ where $n$ is the dimensionality of the solution space. For conciseness and consistency, we report the results with respect to the normalized bandwidth $h_{0}=\frac{h}{W_{\mathrm{domain}}}$ so that $h_{0}\in(0,1]$, where $W_{\mathrm{domain}}$ is the feature space bound for the particular domain.
The experimental results shown in Fig. 4 highlight a peak across all domains around $h_{0}=0.05$. For LP, Deceptive Maze, and Multi-feature LP, the performance decays as the bandwidth increases. However, for Arm Repertoire, DDS-KDE achieves near-optimal coverage for all $h_{0}\geq 0.01$.
@@
Figure 4: Coverage and cross-entropy (CE) after 5,000 iterations of DDS-KDE in all domains for each normalized bandwidth $h_{0}$. We report the mean over 10 trials (3 trials for Deceptive Maze) with error bars showing the standard error of the mean. Higher coverage and lower cross-entropy are better. Highlighted cells are results from the main experiments in Fig. 3.
@@ Appendix C Conversion Formula for Affine Transformations of the Feature Space
$\hat{D}_{h,{\bm{X}}}({\bm{x}};\mathcal{X})=\frac{1}{|\mathcal{X}|h}\sum_{{\bm{x}}^{\prime}\in\mathcal{X}}K\left(\frac{\lVert{\bm{x}}-{\bm{x}}^{\prime}\rVert}{h}\right)$ (8)
@@
$p_{R^{\prime}}(r^{\prime})=\frac{1}{h}K\left(\frac{r^{\prime}}{h}\right)=\frac{1}{h}K\left(\frac{||{\bm{x}}-{\bm{x}}^{\prime}||}{h}\right)$ (9)
@@
$\hat{D}_{h,{\bm{X}}}({\bm{x}};\mathcal{X})=\frac{1}{|\mathcal{X}|h}\sum_{{\bm{x}}^{\prime}\in\mathcal{X}}K\left(\frac{\lVert{\bm{x}}-{\bm{x}}^{\prime}\rVert}{h}\right)=\frac{1}{|\mathcal{X}|}\sum_{{\bm{x}}^{\prime}\in\mathcal{X}}p_{R^{\prime}}(r^{\prime})$ (10)
@@ Appendix D Reservoir Sampling for Maintaining a Representative Buffer
Problem D.1 (Reservoir Sampling (Li,, 1994)).
@@
The population consists of the features that are discovered by DDS throughout its entire execution. Since the buffer ${\mathcal{B}}$ is a simple random sample of all features, the density function estimated on ${\mathcal{B}}$ will be close to the density function of all features. Thus, with reservoir sampling, we are only required to maintain a subset of all discovered features without severely altering the shape of the density estimation of the feature space. In DDS, this is achieved by leveraging algorithm L, the optimal reservoir sampling algorithm (Li,, 1994).
@@ Appendix E Proofs
$\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left|\hat{D}_{h}({\bm{x}};{\mathcal{B}})-\hat{D}_{h}({\bm{x}};{\mathcal{B}}^{\prime})\right|$ $=\sup_{{\bm{x}}\in\mathbb{R}^{m}}\left|\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sum_{{\bm{a}}\in{\mathcal{B}}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}\rVert}{h}\right)-\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sum_{{\bm{a}}^{\prime}\in{\mathcal{B}}^{\prime}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}^{\prime}\rVert}{h}\right)\right|$ (11)
$=\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sup_{{\bm{x}}\in\mathbb{R}^{m}}\left|\sum_{{\bm{a}}\in{\mathcal{B}}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}\rVert}{h}\right)-\sum_{{\bm{a}}^{\prime}\in{\mathcal{B}}^{\prime}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}^{\prime}\rVert}{h}\right)\right|$ (12)
$=\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sup_{{\bm{x}}\in\mathbb{R}^{m}}\left|K\left(\frac{\lVert{\bm{x}}-{\bm{x}}_{i}\rVert}{h}\right)-K\left(\frac{\lVert{\bm{x}}-{\bm{x}}_{j}\rVert}{h}\right)\right|$ (13)
$\leq\frac{1}{|\mathcal{{\mathcal{B}}}|h}$ (14)

### v2

@@ 6.3. Analysis
Figure 3. Coverage and cross-entropy (CE) after 5,000 iterations of each algorithm in all domains. We report the mean over 10 trials, with error bars showing the standard error of the mean. Higher coverage and lower cross-entropy are better.
@@
We attribute the high coverage of DDS-KDE and DDS-CNF on these domains to the continuity of their density estimate, which prevents DDS-KDE from converging prematurely. For example, on LP, our algorithms discover more solutions near the edges of the feature space than CMA-MAE (Appendix G). The passive archive maintained by CMA-MAE converges as all the solutions fall into previously discovered cells (Fontaine et al., 2020). In contrast, the continuity of our density estimate and the online buffer updates facilitate DDS algorithms to achieve slow but continual progress in exploring the feature space (Fig. 3, LP and Arm Repertoire). This is because the shape of the continuous density estimate always changes slightly after each iteration of the algorithm.
For multi-feature LP, we observe that algorithms leveraging continuous representations of the feature space, i.e., DDS and NS, explore the feature space much faster than other algorithms driven by discrete feature space representations, e.g., CMA-MAE (Fig. 3). This is because CMA-MAE is optimizing on a centroidal Voronoi tessellation with 10,000 cells, where each cell has significantly more volume compared to that of a $100\times 100$ grid due to the increased dimensionality of the feature space. Hence, more solutions map to the same cells, making it more difficult to find solutions outside of previously explored cells.
A similar phenomenon was observed in prior work when increasing the dimensionality of the solution space in the LP domain (Fontaine and Nikolaidis, 2023). Increased solution dimensionality more heavily distorts the feature mapping and, similarly, causes most solutions to fall in the same cells of the feature space. Consequently, the LP domain becomes exceptionally challenging for QD algorithms working with a tessellation (like CMA-MAE), as there is insufficient signal to drive the algorithm towards the boundaries of the feature space.
DDS-KDE circumvents this limitation of tessellations with its continuous density estimation of the feature space. While in tesselation-based algorithms, solutions will fall into the same cell, DDS-KDE will retain the solutions in its buffer, which generates signal to drive the search towards the feature space boundaries. Thus, DDS-KDE is more resilient to the distortions of the feature mapping and can better scale with the dimensionality of the feature space.
@@
We attribute the difference between the performance of DDS-KDE and DDS-CNF to the bandwidth parameter in KDE, which allows us to adjust the smoothness of the KDE (2(a)). On the other hand, CNF lacks explicit control over its smoothness. Thus, while we adjust the smoothness of the KDE to improve performance for DDS-KDE (Sec. 5), we can not apply the same techniques to boost the performance of DDS-CNF.
7. Limitations and Future Work
@@
We introduce a new diversity optimization algorithm with two variants, DDS-KDE and DDS-CNF. We provide theoretical insight into the connection between DDS-KDE and novelty search and show that both DDS algorithms excel at discovering diverse solutions.
@@
We envision several extensions to DDS. While we proposed two DDS variants, DDS is a general method that can integrate with a wide variety of distribution fitting techniques. DDS can be combined with parametric estimators like mixture models (McLachlan et al., 2019) and moment matching (Han et al., 2018), as well as with non-parametric estimators like vector quantization (Kohonen, 1990) and generative models (Goodfellow et al., 2020). In general, we believe that DDS will interface with any method that learns a continuous density representation of the feature space.
@@
For DDS-KDE, while we assumed a constant bandwidth, prior work has adapted the bandwidth online (Kristan et al., 2011) and varied the bandwidth based on query location (Terrell and Scott, 1992). These methods could be applied to secure tighter uniform stability bounds for the KDE, making feature space exploration more efficient.
@@
A key limitation of our algorithms is their significant computational cost. Querying the KDE incurs a runtime of $O(|{\mathcal{B}}|)$, where $|{\mathcal{B}}|$ is the buffer size, and fine-tuning the CNF is computationally expensive due to backpropagation, even when accelerated on a GPU. These issues could be mitigated by accelerating KDE computation (O’Brien et al., 2016; Yang et al., 2003) and improving CNF training efficiency (Onken et al., 2021; Huang and Yeh, 2021).
We proposed a generalization of the standard linear projection domain to higher dimensional feature spaces, and demonstrate that DDS exhibits superior performance to existing algorithms. Our results could be further strengthened by evaluating DDS’s performance on more complex domains. For example, we could evaluate the performance of DDS on locomotion tasks, where the solutions are neural network controllers and the task is to control a multi-pedal walker to move efficiently (Sharma et al., 2020; Rakicevic et al., 2021).
@@
Finally, we are excited to see how the underlying insight of DDS — leveraging continuous density representations of the feature space — can improve the exploration power of QD algorithms, especially in high-dimensional feature spaces. Applying DDS to QD will improve its performance on applications such as scenario generation for complex systems (Fontaine et al., 2021), generative design (Gaier et al., 2018), damage recovery in robotics (Cully et al., 2015), reinforcement learning (Batra et al., 2023), and procedural content generation (Gravina et al., 2019).
Acknowledgements.
The authors would like to thank the anonymous referees for their valuable comments and helpful suggestions. This work was supported by the Grant #NSF NRI (#2024949), Grant #NSF GRFP (#DGE-1842487), Grant #NSF CAREER (#2145077), and the NVIDIA Academic Hardware Grant.
@@ Appendix A Hyperparameters
Following prior work (Fontaine et al., 2020), all algorithms run with 15 emitters — emitters (Tjanaka et al., 2023) are objects that generate solutions for QD algorithms. For MAP-Elites (line), the emitter applies the Iso+LineDD operator to random solutions from the archive (Vassiliades and Mouret, 2018). For other algorithms, the emitters are individual instances of CMA-ES. In each experiment, we run the algorithm for 5000 iterations, where each emitter generates $\lambda=36$ solutions per iteration. Following prior work (Fontaine and Nikolaidis, 2023), CMA-ES is configured with the $\mu$ selection rule, basic restart rule, and an initial solution ${\bm{x}}_{0}={\bm{0}}$.
For LP, Arm Repertoire, and Deceptive Maze, MAP-Elites-based algorithms (CMA-MAE, CMA-ME, and MAP-Elites (line)) run on a grid archive of $100\times 100$ following (Fontaine and Nikolaidis, 2023). For Multi-feature LP, since maintaining a grid tessellation would require an exponential number of cells in the number of dimensions, we use a CVT archive with 10,000 cells. Due to randomness in the generation of the CVT cells, we pre-generate the centroids and use the same centroids for all algorithms.
Our DDS algorithms (DDS-KDE and DDS-CNF) maintain a buffer of capacity $b=$10,000$$. DDS-KDE uses a Gaussian kernel with a bandwidth parameter of $h=25.6,10,0.01,2.56$ for Linear Projection, Arm Repertoire, Deceptive Maze, and Multi-feature LP, respectively. The bandwidth is found through empirical fine-tuning (Appendix B).
@@ Appendix B Bandwidth Selection for DDS-KDE
The bandwidth parameter $h$ in KDE controls the smoothness of the estimation (2(a)). While prior work (Silverman, 1986; Rudemo, 1982; Bowman, 1984) has developed bandwidth selection methods for accurate density estimations, our theoretical results show that the optimal bandwidth for DDS-KDE may require an over-smoothing of the KDE (Sec. 5).
To examine the effect of the bandwidth parameter on the performance of DDS-KDE, we perform ablation experiments on LP, Arm Repertoire, and Deceptive Maze. While applying the same hyperparameters as described in Appendix A, we vary the bandwidth $h\in(0,W]$, where $W$ is the bound of the feature space defined for each domain (Sec. 6.2). For instance, in the LP domain, $W_{\mathrm{LP}}=5.12\cdot\frac{n}{2}$ where $n$ is the dimensionality of the solution space. For conciseness and consistency, we report the results in terms of the normalized bandwidth $h_{0}=\frac{h}{W_{\mathrm{domain}}}$ such that $h_{0}\in(0,1]$. $W_{\mathrm{domain}}$ denotes the feature space bound for the particular domain.
We observe a performance peak across all domains around $h_{0}=0.05$ (Fig. 4). For LP, Deceptive Maze, and Multi-feature LP, DDS-KDE’s performance exhibits diminishes as the bandwidth increases for $h_{0}\geq 0.05$. For Arm Repertoire, DDS-KDE achieves near-optimal coverage for all $h_{0}\geq 0.05$.
@@
Figure 4. Coverage and cross-entropy (CE) after 5,000 iterations of DDS-KDE in all domains for each normalized bandwidth $h_{0}$. We report the mean over 10 trials (3 trials for Deceptive Maze) with error bars showing the standard error of the mean. Higher coverage and lower cross-entropy are better. Highlighted cells are results from the main experiments in Fig. 3. The plots show the normalized bandwidth on the $x$-axis
Table 1. Detailed hyperparameter specifications for each algorithm and domain.
DDS-KDE Initial step size $\sigma_{0}$ 1.5 0.5 1.5 1.5
Bandwidth $h$ 25.6 10 0.01 2.56
DDS-CNF Initial step size $\sigma_{0}$ 0.5 1.5 1.5 0.5
NS Initial step size $\sigma_{0}$ 0.5 0.5 0.5 0.5
Nearest neighbors $k$ 100 100 100 100
CMA-MAE Initial step size $\sigma_{0}$ 0.5 0.2 0.5 0.5
Threshold min 0 0 0 0
Learning rate $\alpha$ 0.01 0.01 0.01 0.01
CMA-ME Initial step size $\sigma_{0}$ 0.5 0.2 0.5 0.5
Line variance $\sigma_{\mathrm{line}}$ 0.2 0.2 0.2 0.2
@@ Appendix C Conversion Formula for Affine Transformations of the Feature Space
(8) $\hat{D}_{h,{\bm{X}}}({\bm{x}};\mathcal{X})=\frac{1}{|\mathcal{X}|h}\sum_{{\bm{x}}^{\prime}\in\mathcal{X}}K\left(\frac{\lVert{\bm{x}}-{\bm{x}}^{\prime}\rVert}{h}\right)$
@@
(9) $p_{R^{\prime}}(r^{\prime})=\frac{1}{h}K\left(\frac{r^{\prime}}{h}\right)=\frac{1}{h}K\left(\frac{||{\bm{x}}-{\bm{x}}^{\prime}||}{h}\right)$
@@
(10) $\hat{D}_{h,{\bm{X}}}({\bm{x}};\mathcal{X})=\frac{1}{|\mathcal{X}|h}\sum_{{\bm{x}}^{\prime}\in\mathcal{X}}K\left(\frac{\lVert{\bm{x}}-{\bm{x}}^{\prime}\rVert}{h}\right)=\frac{1}{|\mathcal{X}|}\sum_{{\bm{x}}^{\prime}\in\mathcal{X}}p_{R^{\prime}}(r^{\prime})$
@@ Appendix D Reservoir Sampling for Maintaining a Representative Buffer
Problem D.1 (Reservoir Sampling (Li, 1994)).
@@
The population consists of the features that are discovered by DDS throughout its entire execution. Since the buffer ${\mathcal{B}}$ is a simple random sample of all features, the density function estimated on ${\mathcal{B}}$ will be close to the density function of all features. Thus, with reservoir sampling, we are only required to maintain a subset of all discovered features without severely altering the shape of the density estimation of the feature space. In DDS, this is achieved by leveraging algorithm L, the optimal reservoir sampling algorithm (Li, 1994).
@@ Appendix E Proofs
(11) $\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left|\hat{D}_{h}({\bm{x}};{\mathcal{B}})-\hat{D}_{h}({\bm{x}};{\mathcal{B}}^{\prime})\right|$ $=\sup_{{\bm{x}}\in\mathbb{R}^{m}}\left|\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sum_{{\bm{a}}\in{\mathcal{B}}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}\rVert}{h}\right)-\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sum_{{\bm{a}}^{\prime}\in{\mathcal{B}}^{\prime}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}^{\prime}\rVert}{h}\right)\right|$
(12) $=\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sup_{{\bm{x}}\in\mathbb{R}^{m}}\left|\sum_{{\bm{a}}\in{\mathcal{B}}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}\rVert}{h}\right)-\sum_{{\bm{a}}^{\prime}\in{\mathcal{B}}^{\prime}}K\left(\frac{\lVert{\bm{x}}-{\bm{a}}^{\prime}\rVert}{h}\right)\right|$
(13) $=\frac{1}{|\mathcal{{\mathcal{B}}}|h}\sup_{{\bm{x}}\in\mathbb{R}^{m}}\left|K\left(\frac{\lVert{\bm{x}}-{\bm{x}}_{i}\rVert}{h}\right)-K\left(\frac{\lVert{\bm{x}}-{\bm{x}}_{j}\rVert}{h}\right)\right|$
(14) $\leq\frac{1}{|\mathcal{{\mathcal{B}}}|h}$